OpenStudy (anonymous):

The English Alphabet has 20 letters more than one−fourth the number of letters in the Greek Alphabet. The English Alphabet has 26 letters. Which of the following equations would you use to find the number of letters, g, in the Greek Alphabet?

OpenStudy (nurali):

1/4g + 20 = 26

OpenStudy (goldphenoix):

English Alphabet (26) = 20 + (1/4g) g= Greek Alphabet. So the equation would look like: \[\large \large 20 + \frac{ 1 }{ 4 }g = 26\] Subtract 20 from both side to get 1/4g = 6 Get rid of 1/4 by timing both side by 4/1 to get 24.